OpenStudy (anonymous):

Find the tangent line for y = x^3 at 0. Does the tangent line cross the curve y = x^3 at the point of contact

OpenStudy (anonymous):

i think its 0- what do they mean by point of contact

OpenStudy (anonymous):

point of contact means where the two graphs intersect?

OpenStudy (anonymous):

it seems obvious that they intersect at the point of contact

OpenStudy (turingtest):

\[ y = x^3\rightarrow{dy \over dx}=3x^2\]o is correct the point crosses the curve because it is y^3, if it was y^2 it would not cross it, it would just touch it.

OpenStudy (anonymous):

right X^3

OpenStudy (turingtest):

so along y=0 we will cross this graph

OpenStudy (anonymous):

point of contact means where the two graphs intersect?

OpenStudy (anonymous):

can you show by letting x^3 = 3x^2 then solve for x also?

OpenStudy (anonymous):

so you think thats what they mean by cross? means not just touching?

OpenStudy (turingtest):

if it was y=x^2 it would look likeso y=0 would only TOUCH the graph, not cross it